rosydumbperson:

Math Halp?? A pair of parallel lines is cut by a transversal: A pair of parallel lines cut by a transversal is shown. The inside bottom right angle is labeled 85 degrees. The alternate interior angle is cut by a ray and is labeled as x on one side and 25 degrees on the other. What is the measure of angle x? 50 degrees 55 degrees 60 degrees 80 degrees

dude:

Can you take a ss and post it?

rosydumbperson:

1 attachment
dude:



dude:

 Since they are alternate interior they are the same length

dude:

x+25 = 85

Pixel:

TMW you think you're smart then realize you're till a Nyan Cat

dude:

Its okay pixel :P You're not a "Nyan Cat"

rosydumbperson:

lol wait is 85 the answer? because thats not an option

Pixel:

he isnt done yet

dude:

no you're supposed to solve for x \(x+25=85\)

rosydumbperson:

oh okay

Pixel:

dude never gives answers, he only gives you the path to the answer

dude:

^^^ thank you

rosydumbperson:

ahh thats sad..

Pixel:

how its against the rules to give direct answers

rosydumbperson:

oof.. well i suck at math and i dont understand it

Pixel:

i do too

Pixel:

Thats why i have a Dude

Pixel:

=D

rosydumbperson:

what? im confused

dude:

Solve for x \(x+25=85\) Do you know how to do this?

Pixel:

Dude helps me with math, and i almost always get a 80-100%, most of the time i go too fast

rosydumbperson:

niceeee lol

rosydumbperson:

and no i dont

Pixel:

*face desk*

dude:

Do you want to help on that pixel?

rosydumbperson:

ohhhhh waiiitttttt! i get it. its 60 right?

dude:

Right-o

rosydumbperson:

ahhhh thxxxx

dude:

Sure

rosydumbperson:

ahh thx i passed the test

Pixel:

ggs

Pixel:

dude im lazy can you translate this text for me?

Pixel:

ill post a question

dude:

Make a post yes